YasirH:

Name the number that is three more than one-fifth of one-tenth of one-half of 5,000. A. 503 B. 103 C. 53 D. 108

Timmyspu:

What is a wrong answer?

YasirH:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu What is a wrong answer? \(\color{#0cbb34}{\text{End of Quote}}\) without 1 answer all are the wrong

Timmyspu:

Which one looks to be the least likely to be the answer.

carlosesgirl:

Since, x = three more than one-fifth of one-tenth of one-half of 5,000, Hence, the number is 53

carlosesgirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @YasirH \(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu What is a wrong answer? \(\color{#0cbb34}{\text{End of Quote}}\) without 1 answer all are the wrong \(\color{#0cbb34}{\text{End of Quote}}\) Since, x = three more than one-fifth of one-tenth of one-half of 5,000, Hence, the number is 53

YasirH:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carlosesgirl \(\color{#0cbb34}{\text{Originally Posted by}}\) @YasirH \(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu What is a wrong answer? \(\color{#0cbb34}{\text{End of Quote}}\) without 1 answer all are the wrong \(\color{#0cbb34}{\text{End of Quote}}\) Since, x = three more than one-fifth of one-tenth of one-half of 5,000, Hence, the number is 53 \(\color{#0cbb34}{\text{End of Quote}}\) correct answer

jhonyy9:

Name the number that is three more than one-fifth of one-tenth of one-half of 5,000. x = 3 + (1/5)*(1/10)*(1/2)*5000 try solve this for x x = 3 + (1/5)*(1/10)*2500 x = 3 + (1/5) * 250 x = 3 + 50 x = 53 hope this is more understandably